OpenStudy (anonymous):

i have to write the equation of a line that passes through (4,-2) and is parrallel to 2x-y=6. I know the answer but i am not sure how to go about actually getting the answer. Thanks

OpenStudy (anonymous):

Ok so write the equation in sloep-intercept form.

OpenStudy (anonymous):

ok, its -y=-2x+6

OpenStudy (anonymous):

Which is... y = 2x - 6 right?

OpenStudy (anonymous):

oh ok ya

OpenStudy (anonymous):

So, whenever you have a parallel line, the line you are trying to figure out has the same slope as the line you have.

OpenStudy (anonymous):

ok, so the slope of both lines is 2/1, right?

OpenStudy (anonymous):

Yes, now you know that y = mx + b, so let that be the line you are trying to figure out. So, your line should look like.. y = mx + b We know that m(The Slope) is already 2, so... y = 2x + b Now, just plug in your points to find out b... Plugging (4,-2) into the equation... -2 = 2(4) + b Solve for b and plug it back into y = 2x + b, and your done (BTW, b is the y-intercept) If I helped, please medal me. Thanks!

OpenStudy (anonymous):

Awesome! I just went through and did it like that and got the right answer for the first time!! Thanks you so much i have been trying to figure out how to do this for a while. You were a great help:)

OpenStudy (anonymous):

lol, stupid question. This is my first time on here and i dont know how to "medal" you, how would one do that?

OpenStudy (anonymous):

oh, by the looks of it, i already have. Thanks again :)